OpenStudy (idealist10):

Solve the initial-value problem y'=(y^2-3xy-5x^2)/x^2, y(1)=-1.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

I got y=6x/(1-Cx^6)-x y(1)=6/(1-C)-1=-1 6/(1-C)=0 I can't find C.

OpenStudy (idealist10):

y'=y^2/x^2-3(y/x)-5=u^2-3u-5 u+u'x=u^2-3u-5 u'x=u^2-4u-5 du/(u^2-4u-5)=dx/x A/(u-5)+B/(u+1)=1 A=1/6, B=-1/6 (1/6)ln abs(u-5)-(1/6)ln abs(u+1)=ln abs(x)+C ln(u-5/u+1)^(1/6)=ln abs(x)+C ((u-5)/(u+1))^(1/6)=Cx ((u-5/(u+1)=Cx^6 1-(u-5)/(u+1)=1-Cx^6 (u+1)/(u+1)-(u-5)/(u+1)=1-Cx^6 6/(u+1)=1-Cx^6 u+1=6/(1-Cx^6)

OpenStudy (idealist10):

The answer in the book says y=-x. When I try to find C, I can't because it's undefined.

OpenStudy (anonymous):

Hmm, I'm getting the same result... Even Wolfram can't come up with a particular solution: http://www.wolframalpha.com/input/?i=y%27%3D%28y%5E2-3xy-5x%5E2%29%2Fx%5E2%2C+y%281%29%3D-1 I think the initial value is a typo...

OpenStudy (idealist10):

Yeah, it must be a typo then. Thanks for the help.